# The Prevalence of Resonance Among Young, Close-in Planets

### v1

@@ 1 Introduction
Goldberg & Batygin (2022) suggested that orbital instability may be responsible for the observed intra-system uniformity of planet sizes and regular orbital spacing which is widely known as the ‘peas-in-a-pod’ pattern (Weiss et al., 2018; Millholland et al., 2017; Wang, 2017). See also the simulations by MacDonald et al. (2020), Weiss et al. (2022), and Lammers et al. (2023). In the same vein, the outer Solar System could have been initially resonant (Morbidelli et al., 2007; Thommes et al., 2008; Batygin & Brown, 2010; Nesvorný & Morbidelli, 2012) before being disrupted by planetesimal scattering and ensuing instability, as proposed by the Nice model (Gomes et al., 2005; Tsiganis et al., 2005; Morbidelli et al., 2005).
@@ 2 A Sample of Young Multi-planet Systems
Although we started with the ages reported in the literature, we closely examined the age constraints on a case-by-case basis. The youngest group ($<$ 100 Myr) in our sample plays a pivotal role. We have 7 multi-planet systems whose ages primarily come from their cluster membership. TOI-6109 (Dai et al., 2023b; Vach et al., 2024) is a member of the Alpha Persei cluster / Theia 133 association ($\sim$55-Myr-old, see also $\sim 80$-Myr-old estimate by Boyle & Bouma, 2023). TIC 434398831 (Vach et al., 2024) belongs to Theia 116 ($\sim$48-Myr-old). HD 109833 (Wood et al., 2023) is a member of the Lower Centaurus Crux Association (10 to 16-Myr-old). V1298 Tau (David et al., 2019) belongs to the Taurus-Auriga association ($\sim 23$-Myr-old). We used the updated transit ephemeris of V1298 Tau from Livingston et al. (in prep.). AU Mic (Plavchan et al., 2020) is a pre-main-sequence star with a debris disk in the $\sim$20-Myr-old $\beta$ Pic moving group. TOI-1227 (Mann et al., 2022) is a member of the Lower Centaurus Crux/Musca Group ($\sim$11-Myr-old). It only has one detected transiting exoplanet. However, TTVs analyzed by Almenara et al. (2024) revealed the presence of another planet likely in a 3:2 MMR with the known planet. Although TOI-942 (Zhou et al., 2021; Carleo et al., 2021) is not a cluster member, various age indicators and different groups of researchers agree that the age is $\sim$50-Myr.
@@
To check that any trend we report in this work is not driven by system parameters other than the age, we compared the distribution of stellar parameters ($T_{\rm eff}$, $M_{\star}$, [Fe/H]) and planetary parameters ($R_{p}$, $P_{\rm orb}$) between the different age groups. Planet mass $M_{p}$ is not available for the majority of confirmed planets. We performed Kolmogorov-Smirnov tests implemented in scipy on the various stellar and planetary parameters that are available. To account for measurement uncertainties, we bootstrapped using the uncertainties reported in the literature. The only parameter that is statistically distinct (p-value $<5\%$) between the age groups is the planetary radius. Both the young ($<$100-Myr-old) and the adolescent (0.1-1 Gyr-old) planets have larger radii than the mature planets ($>$1-Gyr-old).
@@ 4 Results and Discussion
The 2:1 resonance is weaker than the 3:2 because the 2:1 involves a larger planet-planet separation, and because the 2:1’s indirect gravitational potential (due to the reflex motion of the star) weakens its disturbing function (Murray-Clay & Chiang 2005). The 2:1 resonance may thus require slower migration rates (lower disk surface densities) to capture planets effectively (Kajtazi et al., 2023; Batygin & Petit, 2023). Additionally, overstable libration (Goldreich & Schlichting, 2014; Deck & Batygin, 2015) may remove planets from the 2:1 resonance. The stronger 3:2 resonance may be the first-encountered resonance that robustly captures migrating planets and tends to survive for longer timescales. Higher-$p$ resonances (4:3, 5:4, 6:5) are progressively less populated because a pair of planets must avoid lower-$p$ resonances to reach them. Another possibility is that higher-$p$ resonances may start to overlap and destabilize (e.g. Hadden & Lithwick, 2018; Petit et al., 2020).
@@ 4.1 Resonances Tend to Congregate in a Given System
To further test this result, we carried out a bootstrap experiment. We generated 100 samples of synthetic multi-planet ($\geq 3$-planet) systems for which there was, by design, no correlation in the occurrence of resonances within the same system. We kept the total number of systems (262) and the planet multiplicity in each system the same as in the observed sample. However, orbital period ratios were randomly drawn from the global period ratio distribution. We computed the same set of resonant statistics for these synthetic samples as we did for the observed sample. The results are shown on the right panel of Fig. 2. The simulated systems have significantly fewer cases where two first-order resonant pairs appeared in the same system (11, down from 61). The other rows in Fig. 2 indicate a similar trend: resonant pairs tend to congregate within systems. See also relevant results reported by Jiang et al. (2020).
@@
A requirement for successful capture into resonance is adiabaticity: disk-driven migration must occur over a timescale longer than the resonant libration period. A lower gaseous disk density, which leads to slower migration, usually favors the formation of resonant chains (Xu & Lai, 2017; Dai et al., 2023a). Conversely, stochastic forces due to disk turbulence can dislodge planets from resonance (e.g. Rein & Papaloizou, 2009; Batygin & Adams, 2017; Wu et al., 2024a). Our finding that the majority of young, close-in planets have near-resonant orbits (Section 4.2) suggests that their parent protoplanetary disks had relatively low gas densities and low turbulence (but see Section 4.5 for a counter-point on disk turbulence). The characteristically low gas-mass fractions of sub-Neptunes, and the need to shed most of the disk gas to allow solid proto-cores to dynamically excite each other and merge into super-Earth cores, also point to gas-poor (but not gas-empty) disks as the birth environments for Kepler planets (Lee & Chiang, 2016). Migration is naturally slow during this last phase of gas disk dispersal.
@@ 4.2 Resonances are Common among Young Planets
Are resonances more common among younger planetary systems? Among the 7 multi-planet systems younger than 100 Myrs, 6/7 (85±$\pm 13$%) show at least one near-resonant (first or second-order) planet pair. The only exception is TOI-942, and even in that case, there is a hint of a possible resonance. The period ratio of TOI-942 b and c is very close to $1.5^{2}$. If an undetected planet were placed at the geometric center, it would be in a 3:2 resonance with both known planets, and the average $\Delta$ could be as low as 0.021. TOI-942 has only been observed in TESS Sector 5, and there are too few transits for a TTV analysis.
@@
The incidence of resonance among adolescent systems (0.1 to 1-Gyr-old) lies between those of the young and mature populations: 6/16 (38$\pm$12%) of adolescent systems contain at least one resonant pair, compared to 6/7 (85$\pm$13%) for the young bin, and 58/255 (22.7$\pm$2.6%) for the mature bin. This trend is echoed when counting pairs: 7/29 ($24.1\pm 7.9$%) of the neighboring adolescent pairs are close to first-order resonances, lower than the 7/10 (70±$\pm 15$%) observed in young pairs, and higher than the 59/402 (14.7±$\pm 1.8$%) observed in mature pairs.
Fig. 6 summarizes the incidence of resonant configurations as a function of age as well as a cumulative distribution of $\Delta$ near first-order resonance. Our results indicate that resonant configurations are practically universal among young $<100$-Myr planetary systems, and that the resonances dissolve over timescales on the order of 100 Myr. These findings directly support the ‘breaking-the-chains’ model (Izidoro et al., 2017, 2021; Goldberg et al., 2022), in which close-in planetary systems are initially assembled into resonant configurations that are gradually disrupted by dynamical instabilities.
@@ 4.5 Resonant Repulsion
$\ \Delta(t)\approx 0.006\left(\frac{Q_{\rm p}^{\prime}}{10}\right)^{-1/3}\left(\frac{M_{\rm p}}{10M_{\oplus}}\right)^{1/3}\left(\frac{R_{\rm p}}{2R_{\oplus}}\right)^{5/3}$ (2)
@@
Among the young systems, TOI-1136 (700-Myr-old; Dai et al., 2023a) and AU Mic (20-Myr-old; Plavchan et al., 2020) are relatively deep in resonance ($\Delta\lesssim 0.1\%$). The inferred reduced tidal quality factors are roughly $10^{4}-10^{5}$. Given their Neptune-like and sub-Neptune radii, such a $Q_{p}^{\prime}$ is not surprising as it is close to that of Uranus (see recent result by Gomes & Correia, 2024). However, HD 109833 (10 to 16-Myr-old; Wood et al., 2023) and TOI-2076 (200-Myr-old; Hedges et al., 2021) are significantly farther from commensurability ($\Delta\sim 1\%$), such that the required $Q^{\prime}_{\rm p}$ is smaller than 10. This demands an implausibly fast dissipation rate, especially considering the planets are sub-Neptune rather than terrestrial in size. Therefore, resonant repulsion from equilibrium eccentricity tides cannot be the full story here.
@@ 6 Summary
  • Young planetary systems ($<$100 Myrs old) seem to be dominated by resonant configurations, with 6/7 (85±$\pm 13$%) having at least one resonant pair (including 1st and 2nd-order). Among neighboring pairs, 7/10 (70±$\pm 15$%) are near 1st-order resonances.

### v2

@@ 1 Introduction
Goldberg & Batygin (2022) suggested that orbital instability may be responsible for the observed intra-system uniformity of planet sizes and regular orbital spacing which is widely known as the ‘peas-in-a-pod’ pattern (Weiss et al., 2018; Millholland et al., 2017; Wang, 2017; Goyal et al., 2023). See also the simulations by MacDonald et al. (2020), Izidoro et al. (2022), Weiss et al. (2022), and Lammers et al. (2023). In the same vein, the outer Solar System could have been initially resonant (Morbidelli et al., 2007; Thommes et al., 2008; Batygin & Brown, 2010; Nesvorný & Morbidelli, 2012) before being disrupted by planetesimal scattering and ensuing instability, as proposed by the Nice model (Gomes et al., 2005; Tsiganis et al., 2005; Morbidelli et al., 2005).
@@ 2 A Sample of Young Multi-planet Systems
Although we started with the ages reported in the literature, we closely examined the age constraints on a case-by-case basis. The youngest group ($<$ 100 Myr) in our sample plays a pivotal role. We have 7 multi-planet systems whose ages primarily come from their cluster membership. TOI-6109 (Dai et al., 2023b; Vach et al., 2024, Dattilo et al. in prep) is a member of the Alpha Persei cluster / Theia 133 association ($\sim$55-Myr-old, see also $\sim 80$-Myr-old estimate by Boyle & Bouma, 2023). TIC 434398831 (Vach et al., 2024) belongs to Theia 116 ($\sim$48-Myr-old). HD 109833 (Wood et al., 2023) is a member of the Lower Centaurus Crux Association (10 to 16-Myr-old). V1298 Tau (David et al., 2019) belongs to the Taurus-Auriga association ($\sim 23$-Myr-old). We used the updated transit ephemeris of V1298 Tau from Livingston et al. (in prep.). AU Mic (Plavchan et al., 2020) is a pre-main-sequence star with a debris disk in the $\sim$20-Myr-old $\beta$ Pic moving group. TOI-1227 (Mann et al., 2022) is a member of the Lower Centaurus Crux/Musca Group ($\sim$11-Myr-old). It only has one detected transiting exoplanet. However, TTVs analyzed by Almenara et al. (2024) revealed the presence of another planet likely in a 3:2 MMR with the known planet. Although TOI-942 (Zhou et al., 2021; Carleo et al., 2021) is not a cluster member, various age indicators and different groups of researchers agree that the age is $\sim$50-Myr.
@@
To check that any trend we report in this work is not driven by system parameters other than the age, we compared the distribution of stellar parameters ($T_{\rm eff}$, $M_{\star}$, [Fe/H]) and planetary parameters ($R_{p}$, $P_{\rm orb}$) between the different age groups. Planet mass $M_{p}$ is not available for the majority of confirmed planets. We performed both Anderson-Darling and Kolmogorov-Smirnov tests implemented in scipy on the various stellar and planetary parameters that are available. To account for measurement uncertainties, we bootstrapped using the uncertainties reported in the literature. The only parameter that is statistically distinct (p-value $<5\%$) between the age groups is the planetary radius. Both the young ($<$100-Myr-old) and the adolescent (0.1-1 Gyr-old) planets have larger radii than the mature planets ($>$1-Gyr-old).
@@ 4 Results and Discussion
The 2:1 resonance is weaker than the 3:2 because the 2:1 involves a larger planet-planet separation, and because the 2:1’s indirect gravitational potential (due to the reflex motion of the star) weakens its disturbing function (Murray-Clay & Chiang 2005). The 2:1 resonance may thus require slower migration rates (lower disk surface densities) to capture planets effectively (Kajtazi et al., 2023; Batygin & Petit, 2023). Additionally, overstable libration (Goldreich & Schlichting, 2014; Deck & Batygin, 2015) may remove planets from the 2:1 resonance. The stronger 3:2 resonance may be the first-encountered resonance that robustly captures migrating planets and tends to survive for longer timescales. Higher-$p$ resonances (4:3, 5:4, 6:5) are progressively less populated because a pair of planets must avoid lower-$p$ resonances to reach them (see Fig. 2). Another possibility is that higher-$p$ resonances may start to overlap and destabilize (e.g. Hadden & Lithwick, 2018; Petit_instability,Rath).
Figure 2: The fractions of observed neighboring pairs of planets across all ages near various first and second-order resonances. Resonances with smaller period ratios (4:3, 5:4, 6:5) are progressively less populated because a pair of planets must avoid all earlier resonances to reach them. Another possibility is that these resonances may start to overlap and destabilize (e.g. Rath et al., 2022).
@@ 4.1 Resonances Tend to Congregate in a Given System
To further test this result, we carried out a bootstrap experiment. We generated 100 samples of synthetic multi-planet ($\geq 3$-planet) systems for which there was, by design, no correlation in the occurrence of resonances within the same system. We kept the total number of systems (262) and the planet multiplicity in each system the same as in the observed sample. However, orbital period ratios were randomly drawn from the global period ratio distribution. We computed the same set of resonant statistics for these synthetic samples as we did for the observed sample. The results are shown on the right panel of Fig. 3. The simulated systems have significantly fewer cases where two first-order resonant pairs appeared in the same system (11, down from 61). We performed a contingency table test as implemented in scipy.chi2_contingency. For systems with an existing first-order resonant pair, it is indeed much more likely to find another resonant pair with a p-value of order 10-7. For systems with an existing second-order resonant pair, it is also more likely to see another resonant pair, although the statistical significance is weaker (p-value = 0.048). This is perhaps because second-order resonance is intrinsically rare: there is a much smaller sample with which we can perform this test (second row of Fig . 3). In other words, resonant pairs tend to congregate within systems. See also relevant results reported by Jiang et al. (2020).
@@
One requirement for successful capture into resonance is adiabaticity: disk-driven migration must occur over a timescale longer than the resonant libration period. Our finding that the majority of young, close-in planets have near-resonant orbits (Section 4.2) may favor slower disk migration and lower disk turbulence. Stochastic forces due to disk turbulence can dislodge planets from resonance (e.g. Rein & Papaloizou, 2009; Batygin & Adams, 2017; Wu et al., 2024a). The rate of disk migration depends on several factors such as disk surface density, radial profile of the disk, viscosity among those. Lee & Chiang (2016) suggested the characteristically low gas-mass fractions of sub-Neptunes, and the need to shed most of the disk gas to allow solid proto-cores to dynamically excite each other and merge into super-Earth cores, also point to gas-poor (but not gas-empty) disks as the birth environments for Kepler planets. Disk migration is naturally slow during this last phase of gas disk dispersal.
@@ 4.2 Resonances are Common among Young Planets
Are resonances more common among younger planetary systems? Among the 7 multi-planet systems younger than 100 Myrs, 6/7 (86±$\pm 13$%) show at least one near-resonant (first or second-order) planet pair. The only exception is TOI-942, and even in that case, there is a hint of a possible resonance. The period ratio of TOI-942 b and c is very close to $1.5^{2}$. If an undetected planet were placed at the geometric center, it would be in a 3:2 resonance with both known planets, and the average $\Delta$ could be as low as 0.021. TOI-942 has only been observed in TESS Sector 5, and there are too few transits for a TTV analysis. We discuss some caveats due to selection bias in Section 5.
@@
The incidence of resonance among adolescent systems (0.1 to 1-Gyr-old) lies between those of the young and mature populations: 6/16 (38$\pm$12%) of adolescent systems contain at least one resonant pair, compared to 6/7 (86$\pm$13%) for the young bin, and 58/255 (22.7$\pm$2.6%) for the mature bin. This trend is echoed when counting pairs: 7/29 ($24.1\pm 7.9$%) of the neighboring adolescent pairs are close to first-order resonances, lower than the 7/10 (70±$\pm 15$%) observed in young pairs, and higher than the 59/402 (14.7±$\pm 1.8$%) observed in mature pairs.
Fig. 7 summarizes the incidence of resonant configurations as a function of age as well as a cumulative distribution of $\Delta$ near first-order resonance. Our results indicate that resonant configurations are practically universal among young $<100$-Myr planetary systems, and that the resonances dissolve over timescales on the order of 100 Myr. An Anderson-Darling test using anderson_ksamp in scipy suggests that the $\Delta$ distribution of young $<100$-Myr planetary systems is markedly distinct from the mature sample ($>1$-Gyr) with a p value of 0.0025. These findings directly support the ‘breaking-the-chains’ model (Izidoro et al., 2017, 2021; Goldberg & Batygin, 2022), in which close-in planetary systems are initially assembled into resonant configurations that are gradually disrupted by dynamical instabilities.
@@ 4.5 Resonant Repulsion
$\ \Delta(t)\approx 0.006\left(\frac{Q_{\rm p}^{\prime}}{100}\right)^{-1/3}\left(\frac{M_{\rm p}}{10M_{\oplus}}\right)^{1/3}\left(\frac{R_{\rm p}}{2R_{\oplus}}\right)^{5/3}$ (2)
@@
Among the young systems, TOI-1136 (700-Myr-old; Dai et al., 2023a) and AU Mic (20-Myr-old; Plavchan et al., 2020) are relatively deep in resonance ($\Delta\lesssim 0.1\%$). The inferred reduced tidal quality factors are roughly $10^{3}-10^{4}$. Given their Neptune-like and sub-Neptune radii, such a $Q_{p}^{\prime}$ is not surprising as it is close to that of Uranus (see recent result by Gomes & Correia, 2024). However, HD 109833 (10 to 16-Myr-old; Wood et al., 2023) and TOI-2076 (200-Myr-old; Hedges et al., 2021) are significantly farther from commensurability ($\Delta\sim 1\%$), such that the required $Q^{\prime}_{\rm p}$ is smaller than 1. This demands an implausibly fast dissipation rate, especially considering the planets are sub-Neptune rather than terrestrial in size. Therefore, resonant repulsion from equilibrium eccentricity tides cannot be the full story here.
@@ 6 Summary
  • Young planetary systems ($<$100 Myrs old) seem to be dominated by resonant configurations, with 6/7 (86±$\pm 13$%) having at least one resonant pair (including 1st and 2nd-order). Among neighboring pairs, 7/10 (70±$\pm 15$%) are near 1st-order resonances. This prevalence of resonance among young planetary systems directly supports the ‘break-the-chains’ models (Izidoro et al., 2017, 2022; Goldberg & Batygin, 2022).